What is the total consumption spending Equation if the saving function is S =-4000 + 0,4Y and autonomous investment spending is 8000


1
Expert's answer
2021-10-06T09:46:24-0400

"S =-4000 + 0.4Y\\
Y =C + S
\\Y =C + (-4000 + 0.4Y)
\\C = 4000 - 0.4Y + Y
\\C = 4000 + 0.6Y"

"C = 4000 + 0.6 \times 8000"

"C = 8800"

"=4000 + 0.6Y"